OpenStudy (anonymous):

find the equation of the tangent line to y=x^2+3x from point 0,-1 show your work

OpenStudy (amistre64):

what y'?

OpenStudy (amistre64):

the slope of the tangent line is equal to the derivative of "y"

OpenStudy (amistre64):

once we know the slope, the rest is natural

OpenStudy (anonymous):

that's all

OpenStudy (anonymous):

there's nothing more

OpenStudy (amistre64):

then we should find the derivative of y.

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

could you teach me how?

OpenStudy (amistre64):

practice makes perfect; the question suggests that you should be familiar with the derivative in some way; i can go thru some simple shortcuts for you

OpenStudy (amistre64):

but to teach you derivatives is beyond the scope of this environment

OpenStudy (anonymous):

alright

OpenStudy (amistre64):

y=x^2+3x the derivative of a sum is the sum of the derivatives y' = x^2' + 3x'

OpenStudy (amistre64):

the derivative of x^n = nx^{n-1} for example: x^5 = 5x^4

OpenStudy (amistre64):

from this, what would you deduce is the derivative of x^2?

OpenStudy (anonymous):

what you just explained made no sense to me

OpenStudy (anonymous):

but don't worry about it

OpenStudy (anonymous):

im not at the level math u r

OpenStudy (amistre64):

you taking calculus yet?

OpenStudy (anonymous):

nope

OpenStudy (amistre64):

thats where you learn about derivatives :)

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

at any rate; if you knew the derivative got us to 2x+3; this would tell us the equation of the slope at any given point on our line

OpenStudy (amistre64):

at x=0 this slope = 3

OpenStudy (amistre64):

given a slope and a point we can construct a line: y = mx -mPx + Py y = 3x -3(0) +-1 y = 3x -1 looks to be it

OpenStudy (amistre64):

good luck with it :)

OpenStudy (anonymous):

ok, thanks